# An Adaptive Dual-Domain Prediction Strategy based on Second-order Derivatives for Dynamic Multi-Objective Optimization

### v1

@@ Abstract
This paper addresses the problem of dynamic multi-objective optimization problems (DMOPs), by demonstrating new approaches to change prediction strategies within an evolutionary algorithm paradigm. Because the objectives of such problems change over time, the Pareto optimal set (PS) and Pareto optimal front (PF) are also dynamic. To accurately track the changing PS and PF in the decision and objective spaces, we propose a novel adaptive prediction strategy, which utilizes the concept of second-order derivatives adaptively in different domains. Firstly, we propose a dual-domain method, which takes into account changes in both the PS and the PF simultaneously. An adaptive strategy is adopted to self-adjust the proportion of the search space. Secondly, a second-order derivative prediction strategy is proposed to predicatively re-initialize the population. We compare the performance of the proposed algorithm against four other state-of-the-art algorithms from the literature, using DMOPs benchmark problems. Experimental results show that the proposed method outperforms the other algorithms on most of the test problems.
@@ I Introduction
Numerous practical problems can be formulated as dynamic multi-objective problems (DMOPs) [1]. In simple terms, a ”dynamic environment” means that the problem is constantly changing over time. For example, imagine you are trying to optimize a delivery route, but the traffic conditions, weather, and customer demands keep changing. This makes finding the best solution more challenging compared to static problems, where conditions remain constant. As a result, the convergence process for DMOPs becomes more complex than that of static multi-objective problems. Without loss of generality, a DMOP can be formulated mathematically as:
@@
$\begin{array}[]{l}\min\;F({\bm{x},t})=\{f_{1}(\bm{x},t),\cdots,f_{m}(\bm{x},t)\}^{T}\\ \text{subject to}\quad\bm{x}\in\displaystyle\prod_{i=1}^{n}\left[{a}_{i},{b}_{i}\right]\\ \end{array}$ (1)
@@
where $F(\bm{x},t)$ is the vector of objective functions with $m$ time-varying objective functions $f_{j}(\bm{x},t)$, $j=1,2,\ldots,m$. $\bm{x}$ is an $n$-dimensional decision variable, $t=1,2,3,\ldots$ $T$. $T$ denotes the time scale, and $a_{i},b_{i}$ are constants in the range of (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) for all $i=1,2,\ldots,n$.
@@
Multi-objective evolutionary algorithms (MOEAs) are recognized as a powerful technique for solving multi-objective optimization problems (MOPs) [2, 3]. There is a particular class of MOPs [2, 4, 5] which becomes important in a variety of science and engineering applications. These MOPs are challenging, in having mutually conflicting objective functions, so that improving the performance of one function often comes at the expense of the other functions. It is impossible to optimize all such functions simultaneously. However, we can obtain a set of equilibrium solutions which is called the Pareto-optimal Set (PS) in decision variable space. This corresponds to a Pareto-optimal Front (PF) in objective function space [6, 7]. However, in certain kinds of real-world applications, the parameters of MOPs may change over time. This causes the optimal PF to change over time. These problems with changing PFs are usually called Dynamic MOPs (DMOPs) [1]. DMOPs can be especially challenging to solve. This is partly because the objective functions are continually changing over time, and also because multiple, mutually contradictory objective functions need to be optimized at each time step. A critical challenge in solving DMOPs is that of tracking the trajectories of changeable PS and PF in various environments.
@@
Dynamic multi-objective optimization is widely applied in dynamic scheduling [8, 9], planning [10, 11, 12, 13], resource allocation [14], and machine learning [15]. It can be difficult to adapt conventional, static MOEAs to problems where the environment changes dynamically. To address this, researchers have developed strategies to improve the convergence of solutions to DMOPs, including the diversity-based [16, 17], memory-based [18, 19, 20, 21], and change prediction-based [22, 23, 24] methods. Diversity-based methods include mutating or re-initializing populations based on the severity of environmental changes [25]. Memory-based methods reuse the past information by storing representative individuals to conduct future searches[18]. Change prediction-based strategies use past search experiences to guide the evolutionary process towards future PS or PF directions, and have gained significant attention in the research community.
@@
However, the methods above encounter several challenges when addressing certain real-world DMOPs, characterized by three principal issues. Firstly, the absence of precise guidance strategies might negatively impact subsequent optimization efforts, as they fail to provide clear directions for the algorithm to follow when changes occur. Secondly, approaches that utilize memory to recall past optimal solutions often struggle to swiftly adapt to new environmental conditions, resulting in outdated or irrelevant solutions being recalled. Thirdly, previous change prediction strategies typically focus solely on the decision space, neglecting the characteristics of the objective space. This one-sided focus can lead to inaccurate convergence and suboptimal performance, as the objective space dynamics contain important additional information for effective optimization.
@@
In this paper, we consider DMOPs with time-varying objective functions. To address these problems, we propose a new dynamic multi-objective prediction strategy that incorporates a new adaptive multi-view strategy and a new second-order derivative-based model. The proposed strategy can adaptively model in both the decision and objective spaces, and use these models to predict subsequent optima, thereby responding more rapidly and accurately. These two key contributions can be summarized as follows.
  • An adaptive multi-view prediction method is proposed to predict the changes of DMOPs from the dual domains of decision and objective spaces. This method enables targeted predictions based on the varying trends within these spaces, and dynamically adjusts the weights of predictions in the dual spaces according to the proportion of non-dominated solutions (NDS) [2]. This enhances the effectiveness of evolutionary search in solving DMOPs. This holistic approach aids in fine-tuning the evolutionary strategies, aligns solutions closely with real-time changes, and significantly improves convergence towards optimal solutions under time-varying conditions.
@@
  • A prediction strategy based on the Second-order Derivative method is proposed to reinitialize the new population when a change occurs. This strategy first utilizes an online k-means approach to cluster individuals, which can effectively capture changes in the population distribution, continuously updating the cluster centroids. Secondly, it leverages Second-order Derivative calculations derived from the historical populations’ clustering information to predict the Second-order Derivative of the PF at different time steps. By analyzing the dynamics of past populations, this method accurately forecasts the Second-order Derivative of the PF movement, enabling proactive adjustments to the population. This predictive approach ensures that evolutionary algorithms adapt effectively to evolving objectives, maintaining operational efficiency, and achieving optimal solutions in dynamic optimization environments.
@@
The remainder of this paper is organized as follows. Section II reviews previous literature on DMOPs. Section III presents the framework of the Second-order Derivative method and the adaptive dual-domain prediction strategy. Section IV presents experiments to evaluate the performance of our method and discusses the results. Section V provides concluding remarks and suggestions for future work.
@@ II Related Work
In this section, we first introduce the characteristics of dynamic multi-objective problems. Then we describe existing dynamic multi-objective evolutionary algorithms which utilise prediction methods. Finally, we present the motivation behind our proposed evolutionary search method for solving DMOPs.
@@ II-B Related Dynamic Multi-Objective Evolutionary Algorithms
In recent years, there has been a growing body of research literature on using evolutionary algorithms (EA) to solve DMOPs [1, 22, 23, 24, 27]. Generally, the workflow of existing prediction methods includes two parts: change detection [28] and response strategy. After an environmental change, the first step is to determine if the characteristics of the problem have changed. If they have, a change response strategy, such as prediction stategy, is adopted; otherwise, a static optimization strategy is used.
The DMOEAs are primarily categorized into three types: diversity maintaining methods [25, 29, 30], memory-based methods [20, 21] and prediction methods [22, 23, 24, 27]. These strategies have been developed over recent years and have proven both theoretically and experimentally effective in addressing a wide range of DMOPs.
Diversity maintaining methods [25, 29, 30] are designed to enhance and preserve the population’s diversity upon detecting changes. Liu et al. [25] developed a diversity introduction algorithm that adaptively adjusts individuals to new environmental conditions. Chen et al. [31] have explored an evolutionary method that preserves the diversity of a collection of individuals by treating it as an additional objective in dynamic environments. Peng et al. [29] introduced a technique that uses historical data to guide the search process and increase diversity. These methods initiate individuals in the new environment through random initialization or mutation for mild changes, or through complete re-initialization for more significant changes.
@@ III-C Second-order Derivative Prediction Strategy
${A^{t}}={C^{t{\rm{+}}1}}-2\times{C^{t}}+{C^{t{\rm{-}}1}}$ (7)
@@ IV-B Performance Metrics
1) Invert Generation Distance (IGD) [44]: The IGD metric evaluates both the diversity and convergence of the obtained PF. It is calculated as follows:
@@ IV-D1 MIGD and MHV Values of Final Populations
The statistical results of MIGD and MHV obtained by five algorithms over 20 independent runs in dynamic environment problems, are shown in Table III and Table IV. The mean and variance values for the 20 runs are presented. The best results for each benchmark problem are shown in bold. In addition, we conduct the Wilcoxon rank sum test [49] with a 5% significance level on the outcomes, to demonstrate the performance disparities between the baseline and the proposed algorithm. The symbols ”+”, ”-”and ”=” denote our proposed ADPS-MOEA/D being statistically significantly better, worse, or equivalent to the compared algorithm, respectively. We present a summary of the significance test results along with the MIGD and MHV scores for 14 DMOP benchmarks across four dynamic settings (amounting to 56 DF instances in total) in the bottom row of each table. The findings suggest that the proposed ADPS-MOEA/D outperforms the comparison algorithms using the same optimizer.
To explore whether the severity of dynamic changes influence to the comparative performance, we set four groups of various parameters, donated as C1 and C2, etc. respectively, as described above. The comparative results are shown in Table III and Table IV. From the numerical results we can see that all the five algorithms are sensitive to the severity of change. In addition, the large values of ${n_{t}}$ cause the MIGD and MHV values to become smaller for most of the test problems. Therefore, all the methods demonstrate some ability to make good predictions of future promising solutions. However, for most test problems, the MIGD of ADPS is smaller than those of its counterparts, which shows the advantages of the proposed method. ADPS has better solution distribution on DF1, DF2, DF3, DF7, DF8 and DF10. as MHV are greater than the other four methods with various ${n_{t}}$. This indicates that the obtained solution set is uniform and well-distributed. From the metrics of DF3, it can be observed that the results of the PPS and SVR-MOEA/D methods are slightly worse, while the MV-MOEA/D and ADPS-MOEA/D perform similarly in the MIGD metric.
@@
    TABLE IV: Mean and Standard Deviations of MHV Values Obtained by the Compared Algorithms with Four Dynamic Configurations
    Problem SVR-MOEA/D PPS MV-MOEA/D IGP-DMOEA ADPS-MOEA/D
    DF1 C1 1.2249e+0 (2.23e-2) + 9.3251e-1 (2.78e-2) + 1.5935e+0 (3.57e-2) + 1.5821e+0 (3.09e-2) + 1.7269e+0 (8.11e-3)
    C2 1.6024e+0 (2.46e-2) + 1.1119e+0 (5.16e-2) + 1.6904e+0 (1.94e-2) + 1.7239e+0 (1.21e-2) + 1.7506e+0 (6.57e-3)
    C3 1.3275e+0 (3.93e-2) + 1.2487e+0 (6.87e-2) + 1.7035e+0 (1.81e-2) + 1.6179e+0 (3.16e-2) + 1.7480e+0 (7.01e-3)
    C4 1.6640e+0(1.72e-2) + 1.4809e+0 (6.59e-2) + 1.7512e+0 (7.55e-3) + 1.7601e+0 (1.33e-2) = 1.7648e+0 (4.47e-3)
    DF2 C1 1.5424e+0 (2.46e-2) + 1.3479e+0 (5.69e-2) + 1.7502e+0 (2.26e-2) + 1.6785e+0 (2.72e-2) + 1.9181e+0 (1.57e-2)
    C2 1.7885e+0 (1.86e-2) + 1.6066e+0 (6.20e-2) + 1.8435e+0 (1.47e-2) + 1.8687e+0 (8.07e-3) + 1.9555e+0 (2.84e-3)
    C3 1.6655e+0 (2.08e-2) + 1.4360e+0 (6.79e-2) + 1.8319e+0 (1.04e-2) + 1.6313e+0 (1.93e-2) + 1.9429e+0 (4.86e-3)
    C4 1.8226e+0 (1.04e-2) + 1.7463e+0 (2.99e-2) + 1.8948e+0 (1.30e-2) + 1.8589e+0 (9.89e-3) + 1.9575e+0 (2.83e-3)
    DF3 C1 9.1720e-1 (5.95e-2) + 1.5887e+0 (1.93e-2) + 1.5347e+0 (2.37e-2) + 6.7159e-1 (4.46e-2) + 1.7001e+0 (2.66e-2)
    C2 1.1215e+0 (2.86e-2) + 1.6900e+0 (8.46e-3) + 1.6153e+0 (3.18e-2) + 1.4480e+0 (2.12e-1) + 1.7264e+0 (7.66e-3)
    C3 9.2387e-1 (4.27e-2) + 1.6737e+0 (1.08e-2) + 1.6277e+0 (3.04e-2) + 7.7979e-1 (9.84e-2) + 1.7289e+0 (5.54e-3)
    C4 1.1558e+0 (1.35e-2) + 1.7236e+0 (6.26e-3) + 1.7021e+0 (1.30e-2) + 1.6701e+0 (1.23e-1) + 1.7430e+0 (2.32e-3)
    DF4 C1 8.3960e+0 (1.59e-2) + 8.5476e+0 (2.42e-2) + 8.6003e+0 (1.55e-2) = 7.9157e+0 (1.89e-1) + 8.6019e+0 (2.00e-2)
    C2 8.5510e+0 (3.76e-3) + 8.6302e+0 (9.68e-3) + 8.6407e+0 (6.30e-3) + 8.2382e+0 (4.62e-2) + 8.6552e+0 (6.37e-3)
    C3 8.2707e+0 (1.19e-2) + 8.5147e+0 (8.71e-3) + 8.5229e+0 (6.98e-3) + 7.8118e+0 (1.34e-1) + 8.5363e+0 (7.27e-3)
    C4 8.4717e+0 (4.78e-3) + 8.5428e+0 (2.85e-3) + 8.5393e+0 (1.98e-3) + 8.2740e+0 (1.41e-1) + 8.5486e+0 (1.57e-3)
    DF5 C1 7.3958e-1 (3.68e-2) + 6.6697e-1 (9.69e-2) + 1.7074e+0 (2.69e-2) - 7.7423e-1 (6.37e-2) + 1.6164e+0 (5.35e-2)
    C2 1.6608e+0 (2.84e-2) + 1.0851e+0 (1.15e-1) + 1.7755e+0 (1.53e-2) - 1.4293e+0 (5.92e-2) + 1.6999e+0 (6.63e-2)
    C3 1.3619e+0 (1.88e-2) + 1.1471e+0 (1.76e-1) + 1.7907e+0 (8.15e-3) - 8.0926e-1 (9.00e-2) + 1.6452e+0 (9.02e-2)
    C4 1.8141e+0 (4.31e-3) - 1.5804e+0 (5.68e-2) + 1.7243e+0 (1.41e-2) + 1.6114e+0 (5.77e-2) + 1.7414e+0 (2.72e-2)
    DF6 C1 6.6784e-1 (3.61e-2) + 2.0493e-1 (5.59e-2) + 1.5011e+0 (3.58e-2) - 8.4601e-2 (2.33e-2) + 6.7370e-1 (1.09e-1)
    C2 7.3571e-1 (2.90e-2) + 4.9460e-1 (8.03e-2) + 1.4960e+0 (2.76e-2) - 2.1058e-1 (3.75e-2) + 8.6256e-1 (1.03e-1)
    C3 5.5811e-1 (5.01e-2) + 5.4489e-1 (1.14e-1) + 1.4475e+0 (3.97e-2) - 9.8897e-2 (2.65e-2) + 9.3454e-1 (9.41e-2)
    C4 9.7946e-1 (8.91e-2) + 8.5155e-1 (1.20e-1) + 1.4639e+0 (3.34e-2) - 2.7102e-1 (6.54e-2) + 1.0348e+0 (1.48e-1)
    DF7 C1 5.0769e+0 (1.81e-2) + 5.1488e+0 (8.41e-2) + 5.2353e+0 (6.32e-2) + 5.1405e+0 (6.32e-2) + 5.3913e+0 (9.69e-4)
    C2 5.2972e+0 (4.63e-3) + 5.3098e+0 (3.70e-2) + 5.3538e+0 (3.09e-2) + 5.3880e+0 (2.08e-3) + 5.3937e+0 (4.38e-4)
    C3 5.0474e+0 (6.89e-3) + 5.0593e+0 (2.97e-2) + 5.0758e+0 (1.26e-2) + 4.8720e+0 (5.89e-2) + 5.1215e+0 (1.19e-3)
    C4 5.0979e+0 (4.03e-3) + 5.0852e+0 (1.66e-2) + 5.0978e+0 (1.26e-2) + 5.1076e+0 (4.73e-3) + 5.1236e+0 (3.80e-4)
    DF8 C1 1.8538e+0 (1.99e-3) + 1.8618e+0 (7.90e-5) + 1.8622e+0 (1.02e-4) + 1.8626e+0 (3.73e-5) - 1.8624e+0 (8.36e-5)
    C2 1.8576e+0 (1.10e-3) + 1.8620e+0 (4.05e-5) + 1.8624e+0 (7.15e-5) + 1.8626e+0 (2.39e-5) - 1.8625e+0 (7.07e-5)
    C3 1.8179e+0 (2.00e-3) + 1.8274e+0 (1.17e-4) + 1.8286e+0 (1.38e-4) + 1.8280e+0 (3.67e-5) + 1.8290e+0 (1.45e-4)
    C4 1.8232e+0 (6.05e-4) + 1.8279e+0 (8.72e-5) + 1.8289e+0 (1.65e-4) + 1.8280e+0 (7.85e-5) + 1.8294e+0 (2.14e-4)
    DF9 C1 8.7947e-01 (3.14e-2) + 4.2899e-1 (9.60e-2) + 1.7003e+0 (1.36e-2) - 5.3736e-1 (3.89e-2) + 1.5965e+0 (2.85e-2)
    C2 1.4011e+0 (3.22e-2) + 7.7549e-1 (9.41e-2) + 1.7081e+0 (7.87e-3) = 1.1949e+0 (8.74e-2) + 1.7089e+0 (1.89e-2)
    C3 8.9342e-1 (2.65e-2) + 5.9913e-1 (2.02e-1) + 1.7507e+0 (5.96e-3) - 6.6295e-1 (4.83e-2) + 1.7126e+0 (1.35e-2)
    C4 1.4261e+0 (2.51e-2) + 1.2445e+0 (1.06e-1) + 1.7830e+0 (1.92e-3) - 1.4532e+0 (2.86e-2) + 1.7626e+0 (7.68e-3)
    DF10 C1 1.8197e+0 (2.38e-2) + 1.8067e+0 (1.58e-1) + 2.7677e+0 (3.53e-2) - 1.2560e+0 (1.62e-1) + 2.6512e+0 (7.59e-2)
    C2 2.6139e+0 (2.15e-2) + 2.1690e+0 (1.12e-1) + 2.8028e+0 (1.77e-2) - 1.9345e+0 (2.26e-1) + 2.6999e+0 (1.16e-1)
    C3 1.8052e+0 (3.29e-2) + 1.9946e+0 (1.12e-1) + 2.7283e+0 (2.26e-2) - 1.2669e+0 (3.48e-1) + 2.7092e+0 (3.16e-2)
    C4 2.5564e+0 (1.01e-2) + 2.2459e+0 (7.97e-2) + 2.7325e+0 (2.35e-2) + 1.4909e+0 (4.02e-1) + 2.7470e+0 (3.30e-2)
    DF11 C1 5.3302e+0 (4.68e-3) + 5.3964e+0 (1.49e-2) + 5.4334e+0 (1.20e-2) + 5.3936e+0 (2.79e-2) + 5.4561e+0 (7.25e-3)
    C2 5.4542e+0 (2.04e-3) + 5.4316e+0 (6.95e-3) + 5.4565e+0 (8.67e-3) + 5.4623e+0 (1.49e-2) + 5.4726e+0 (6.10e-3)
    C3 5.4320e+0 (6.10e-3) + 5.4975e+0 (9.27e-3) + 5.5241e+0 (8.44e-3) + 5.5226e+0 (2.47e-2) + 5.5529e+0 (9.20e-3)
    C4 5.5229e+0 (3.89e-3) + 5.5261e+0 (7.33e-3) + 5.5421e+0 (6.58e-3) + 5.5687e+0 (1.24e-2) = 5.5675e+0 (6.61e-3)
    DF12 C1 3.9211e+0 (3.77e-2) + 3.7621e+0 (9.78e-2) + 4.1662e+0 (2.43e-2) + 3.4093e+0 (7.40e-2) + 4.2139e+0 (2.32e-2)
    C2 4.1745e+0 (9.34e-3) + 4.0188e+0 (7.32e-2) + 4.2052e+0 (1.46e-2) + 3.8962e+0 (1.42e-1) + 4.2280e+0 (2.33e-2)
    C3 3.8114e+0 (2.43e-2) + 3.7583e+0 (7.96e-2) + 4.1212e+0 (2.95e-2) - 3.3645e+0 (1.08e-1) + 4.0786e+0 (3.14e-2)
    C4 4.0050e+0 (5.96e-3) + 3.9119e+0 (4.50e-2) + 4.1285e+0 (1.91e-2) - 4.0631e+0 (4.85e-2) + 4.0953e+0 (1.48e-2)
    DF13 C1 8.5736e+0 (3.79e-2) + 8.9729e+0 (1.24e-1) + 9.5317e+0 (1.70e-2) - 8.4348e+0 (4.20e-1) + 9.4279e+0 (4.92e-2)
    C2 9.3411e+0 (1.64e-2) + 9.2308e+0 (6.45e-2) + 9.5462e+0 (1.30e-2) - 8.8277e+0 (4.28e-1) + 9.4860e+0 (2.86e-2)
    C3 9.1029e+0 (3.53e-2) + 9.2026e+0 (7.88e-2) + 9.4306e+0 (2.86e-2) + 8.6873e+0 (4.59e-1) + 9.4655e+0 (3.21e-2)
    C4 9.3433e+0 (1.13e-2) + 9.3499e+0 (3.77e-2) + 9.5278e+0 (1.59e-2) - 9.0824e+0 (2.09e-1) + 9.4786e+0 (1.34e-2)
    DF14 C1 1.0964e+0 (3.13e-2) + 8.0885e-1 (2.25e-1) + 1.6096e+0 (2.34e-2) - 8.5444e-1 (1.31e-1) + 1.2458e+0 (1.46e-1)
    C2 1.4678e+0 (2.63e-2) + 1.1147e+0 (1.88e-1) + 1.6013e+0 (3.28e-2) - 1.2844e+0 (1.24e-1) = 1.3529e+0 (9.20e-2)
    C3 1.2089e+0 (2.80e-2) + 1.0844e+0 (2.16e-1) + 1.6873e+0 (1.66e-2) - 9.6710e-1 (8.72e-2) + 1.3740e+0 (1.89e-1)
    C4 1.6016e+0 (1.18e-2) + 1.3708e+0 (9.10e-2) + 1.6670e+0 (4.19e-2) = 1.5903e+0 (8.68e-2) + 1.7456e+0 (4.02e-2)
    +/-/= 55/1/0 56/0/0 32/21/3 51/2/3 ——
    best/all 1/56 0/56 21/56 3/56 31/56
@@
Figure 6: Convergence curves of the averaged IGD (over 20 runs) obtained by ADPS, ADPS-I and ADPS-II on benchmark DMOPs, with dynamic configuration C3 (vertical axis: IGD value; horizontal axis: index of dynamic change). Left hand graph shows benchmark problem DF1. Right hand graph shows benchmark problem DF13.

### v2

# An accelerate Prediction Strategy for Dynamic Multi-Objective Optimization

@@ Abstract
This paper addresses the challenge of dynamic multi-objective optimization problems (DMOPs) by introducing novel approaches for accelerating prediction strategies within the evolutionary algorithm framework. Since the objectives of DMOPs evolve over time, both the Pareto optimal set (PS) and the Pareto optimal front (PF) are dynamic. To effectively track the changes in the PS and PF in both decision and objective spaces, we propose an adaptive prediction strategy that incorporates second-order derivatives to predict and adjust the algorithms search behavior. This strategy enhances the algorithm’s ability to anticipate changes in the environment, allowing for more efficient population re-initialization. We evaluate the performance of the proposed method against four state-of-the-art algorithms using standard DMOPs benchmark problems. Experimental results demonstrate that the proposed approach significantly outperforms the other algorithms across most test problems.
@@ I Introduction
Numerous practical problems can be formulated as dynamic multi-objective problems (DMOPs) [1]. In simple terms, a ”dynamic environment” means that the problem is constantly changing over time. For example, imagine you are trying to optimize a delivery route, but the traffic conditions, weather, and customer demands keep changing. This makes finding the best solution more challenging compared to static problems, because the conditions remain constant. As a result, the convergence process for DMOPs becomes more complex than that of static multi-objective problems. Without loss of generality, a DMOP can be formulated mathematically as:
@@
$\begin{array}[]{l}\min\;F({\bm{x},t})=\{f_{1}(\bm{x},t),\cdots,f_{m}(\bm{x},t)\}^{T}\\ \text{subject to}\quad\bm{x}\in\Omega=\displaystyle\prod_{i=1}^{n}\left[{a}_{i},{b}_{i}\right]\\ \end{array}$ (1)
@@
where $F(\bm{x},t)$ is the vector of objective functions with $m$ time-varying objective functions $f_{j}(\bm{x},t)$, $j=1,2,\ldots,m$. $\bm{x}$ is an $n$-dimensional decision variable, $t=1,2,3,\ldots$ $T$. $T$ denotes the time scale, and $a_{i},b_{i}$ are the bounds of $\bm{x_{i}}$ , and these constants fall within the range (,+)($\infty$,$\infty$)( - ∞ , + ∞ ) for all $i=1,2,\ldots,n$.
@@
Multi-objective evolutionary algorithms (MOEAs) are recognized as a powerful technique for solving multi-objective optimization problems (MOPs) [2, 3]. There is a particular class of MOPs [2, 4, 5] which becomes important in various scientific and engineering applications. These MOPs are challenging, in having mutually conflicting objective functions, meaning that improving the performance of one function often comes at the expense of the others. It is impossible to optimize all such functions simultaneously. However, we can obtain a set of equilibrium solutions which is called the Pareto-optimal Set (PS) in decision variable space, which corresponds to the Pareto-optimal Front (PF) in the objective function space [6, 7]. However, in certain kinds of real-world applications, the parameters of MOPs may change over time, causing the optimal PF to evolve. These problems with changing PFs are usually called DMOPs [1]. DMOPs can be especially challenging to solve. This is partly because the objective functions are continually changing over time, and also because multiple, mutually contradictory objective functions need to be optimized at each time step. A critical challenge in solving DMOPs is that of tracking the trajectories of the changing PS and PF in various environments.
@@
Dynamic multi-objective optimization is widely applied in dynamic scheduling [8, 9, 10], planning [11, 12, 13, 14], resource allocation [15], and machine learning [16]. It can be difficult to adapt conventional, static MOEAs to problems where the environment changes dynamically. To address this, researchers have developed strategies to improve the convergence of solutions to DMOPs, including the diversity-based [17, 18, 19], memory-based [20, 21, 22], and change prediction-based [23, 24, 25] methods. Diversity-based methods include mutating or re-initializing populations based on the severity of environmental changes. Memory-based methods reuse the past information by storing representative individuals to conduct future searches. Change prediction-based strategies use past search experiences to guide the evolutionary process towards future PS or PF directions, and have gained significant attention in the research community.
@@
However, the methods above encounter several challenges when addressing certain real-world DMOPs, characterized by three main issues. First, the absence of precise guidance strategies might negatively impact subsequent optimization efforts, as they fail to provide clear directions for the algorithm to follow when changes occur. Second, approaches that utilize memory to recall past solutions often struggle to swiftly adapt to new environmental conditions, resulting in outdated or irrelevant solutions being recalled. Third, previous change prediction strategies typically focus solely on the decision space, overlooking the characteristics of the objective space. This narrow focus can lead to inaccurate convergence and suboptimal performance, as the objective space dynamics contain important additional information for effective optimization.
@@
In this paper, we consider DMOPs with time-varying objective functions. To address these problems, we propose a novel dynamic multi-objective prediction strategy that incorporates an adaptive multi-view approach and a second-order derivative-based model. The proposed strategy can adaptively model in both the decision and objective spaces, and use these models to predict subsequent optima, thereby responding more rapidly and accurately. These two key contributions can be summarized as follows.
@@
  • A prediction strategy based on the second-order derivative method is proposed to reinitialize the population when a change occurs. First, this strategy applies an online k-means approach to cluster individuals, effectively capturing changes in the population distribution by continuously updating cluster centroids. Second, it leverages second-order derivative calculations from the clustering information of past populations to predict the PF’s movement at different time steps. By analyzing population dynamics over time, this method accurately forecasts changes in the PF’s trajectory, enabling proactive adjustments to the population. This predictive approach ensures that evolutionary algorithms can adapt efficiently to changing objectives, maintain operational effectiveness, and achieve optimal solutions in dynamic environments.
@@
The remainder of this paper is organized as follows. Section II reviews previous literature on DMOPs. Section III presents the framework of the Second-order Derivative method and the adaptive dual-domain prediction strategy. Section IV presents experiments that evaluate the performance of our method and discusses the results. Section V provides concluding remarks and suggestions for future work.
@@ II Related Work
In this section, we first introduce the characteristics of dynamic multi-objective problems. Then we describe existing dynamic multi-objective evolutionary algorithms. Finally, we present the motivation behind our proposed evolutionary search method for solving DMOPs.
@@ II-B Related Dynamic Multi-Objective Evolutionary Algorithms
In recent years, there has been a growing body of research literature on using evolutionary algorithms (EA) to solve DMOPs [27, 28, 29, 30]. Generally, the workflow of existing prediction methods includes two parts: change detection [31] and response strategies. After an environmental change, the first step is to determine if the characteristics of the problem have changed. If they have, a change response strategy, such as prediction strategy, is adopted; otherwise, a static optimization strategy is used.
The DMOEAs are primarily categorized into three types: diversity maintaining methods [19, 17, 18], memory-based methods [22, 21, 20] and prediction methods [32, 33, 24, 25]. These strategies have been developed over recent years and have been shown to be effective both theoretically and experimentally in addressing a wide range of DMOPs.
Diversity maintaining methods [17, 18] are designed to enhance and preserve the population’s diversity upon detecting changes. Liu et al. [34] developed a diversity introduction algorithm that adaptively adjusts individuals to new environmental conditions. Chen et al. [35] have explored an evolutionary method that preserves the diversity of a collection of individuals by treating it as an additional objective in dynamic environments. Peng et al. [36] introduced a technique that uses historical data to guide the search process and increase diversity. These methods initiate individuals in the new environment through random initialization or mutation for mild changes, or through complete re-initialization for more significant changes.
@@ III-C Second-order Derivative Prediction Strategy
${A^{t}}=\frac{C^{t+1}-2C^{t}+C^{t-1}}{\Delta t^{2}}$ (7)
where $\Delta t$ is the time interval between adjacent time steps, and the acceleration $A^{t}$ is the second-order finite difference approximation of position change over time.
@@ IV-B Performance Metrics
1) Invert Generation Distance (IGD) [49]: The IGD metric evaluates both the diversity and convergence of the obtained PF. It is calculated as follows: (wrong)
@@ IV-D1 MIGD and MHV Values of Final Populations
The statistical results of MIGD and MHV obtained by five algorithms over 20 independent runs in dynamic environment problems, are shown in Table III. The mean and variance values for the 20 runs are presented. The best results for each benchmark problem are shown in bold. In addition, we conduct the Wilcoxon rank sum test [54] with a 5% significance level on the outcomes, to demonstrate the performance disparities between the baseline and the proposed algorithm. The symbols ”+”, ”-”and ”=” denote our proposed ADPS-MOEA/D being statistically significantly better, worse, or equivalent to the compared algorithm, respectively. We present a summary of the significance test results along with the MIGD and MHV scores for 14 DMOP benchmarks across four dynamic settings (amounting to 56 DF instances in total) in the bottom row of each table. The findings suggest that the proposed ADPS-MOEA/D outperforms the comparison algorithms using the same optimizer.
To explore whether the severity of dynamic changes influence to the comparative performance, we set four groups of various parameters, donated as C1 and C2, etc. respectively, as described above. The comparative results are shown in Table III. From the numerical results we can see that all the five algorithms are sensitive to the severity of change. In addition, the large values of ${n_{t}}$ cause the MIGD and MHV values to become smaller for most of the test problems. Therefore, all the methods demonstrate some ability to make good predictions of future promising solutions. However, for most test problems, the MIGD of ADPS is smaller than those of its counterparts, which shows the advantages of the proposed method. ADPS has better solution distribution on DF1, DF2, DF3, DF7, DF8 and DF10. as MHV are greater than the other four methods with various ${n_{t}}$. This indicates that the obtained solution set is uniform and well-distributed. From the metrics of DF3, it can be observed that the results of the PPS and SVR-MOEA/D methods are slightly worse, while the MV-MOEA/D and ADPS-MOEA/D perform similarly in the MIGD metric.